# Improving Conditional VAE with approximation using Normalizing Flows

### v1

@@ Abstract
Variational Autoencoders and Generative Adversarial Networks remained the state-of-the-art (SOTA) generative models until 2022. Now they are superseded by diffusion based models. Efforts to improve traditional models have stagnated as a result. In old-school fashion, we explore image generation with conditional Variational Autoencoders (CVAE) to incorporate desired attributes within the images. VAEs are known to produce blurry images with less diversity, we refer a method that solve this issue by leveraging the variance of the gaussian decoder as a learnable parameter during training. Previous works on CVAEs assumed that the conditional distribution of the latent space given the labels is equal to the prior distribution, which is not the case in reality. We show that estimating it using normalizing flows results in better image generation than existing methods by reducing the FID by 5% and increasing log likelihood by 7.7% than the previous case.
@@ 3.1 Experiments
  • Gaussian CVAE: CVAE with Gaussian decoder where the variance of the decoder is kept as unit and not learned. Typically, the vanilla CVAE and latent space is dependent on the labels.
  • $\sigma$-CVAE (non-NF): The scenario where the decoder of the variance is optimal as described and the latent space $z$ is independent of the labels. This sets $p(z|y)=p(z)$ the in objective equation.
@@ 3.2 Results
Gaussian VAE -4218 300.9
$\sigma$- VAE (non-NF) -6223 173.50
$\sigma$- VAE (NF) -6702 164.48
Table 1: Metrics of VAEs under given scenarios. NLL stands for Negative Log Likelihood on the reconstructed data (lower is better).
We compare the negative log likelihood (NLL) and the Fréchet Inception Distance (FID) FID, FID-github-repo score for all three settings. Table 1 shows the comparative metrics for all three scenarios. The NLL is calculated on the test set during the training. FID is calculated post training by randomly generating images using the labels from the test set and comparing them with the original images of the test set. We observe that VAEs with approximated conditionals outperform all the other VAEs in the settings.

### v3

# CS 5100: Foundations of AI Course Project Improving Conditional VAE with approximation using Normalizing Flows

@@ Abstract
Variational Autoencoders and Generative Adversarial Networks remained the state-of-the-art (SOTA) generative models until 2022. Now they are superseded by diffusion based models. Efforts to improve traditional models have stagnated as a result. In old-school fashion, we explore image generation with conditional Variational Autoencoders (CVAE) to incorporate desired attributes within the images. VAEs are known to produce blurry images with less diversity; we refer to a method that solves this issue by leveraging the variance of the gaussian decoder as a learnable parameter during training. Previous works on CVAEs assumed that the conditional distribution of the latent space given the labels is equal to the prior distribution, which is not the case in reality. We show that estimating it using normalizing flows results in better image generation than existing methods by reducing the FID by 4% and increasing log likelihood by 7.6% compared to the previous cases  [^1] The code is publicly available at: https://github.com/Gituhin/Conditional-VAE-NF.
@@ 3.1 Experiments
  • Gaussian CVAE: CVAE with Gaussian decoder where the variance of the decoder is kept as 1 and not learned. Typically, the vanilla CVAE and latent space is dependent on the labels.
  • $\sigma$-CVAE (non-NF): The scenario where the variance of the decoder is optimal as described and the latent space $z$ is independent of the labels. This sets $p(z|y)=p(z)$ the in objective equation.
@@ 3.2 Results
Gaussian VAE -32.95 389.20 389.06
$\sigma$- VAE (NNF) -48.61 107.83 166.07
$\sigma$- VAE (NF) -52.32 107.24 159.13
Table 1: Metrics of VAEs under given scenarios. NLL stands for Negative Log Likelihood on the reconstructed data (lower is better). FID (Recon) is the score on reconstructed images from the test set, whereas FID (Sampled) is the score on randomly sampled images and the test set. (See text for details)
We compare the negative log likelihood (NLL) and the Fréchet Inception Distance (FID) FID, FID-github-repo score for all three settings. Table 1 shows the comparative metrics for all three scenarios. The NLL is calculated on the test set during the training. There are two FID scores, both of which are calculated post training. FID (Recon) or reconstructed is calculated between the reconstructed images test set images and the original test set. FID (Sampled) is calculated between the sampled images from random latent spaces with the attributes from the test set and the original test set images. We observe that CVAEs with approximated conditionals outperform all the other CVAEs in the settings.